Bellingham Company produces a product that requires 10 standard direct labor hours per unit at a standard hourly rate of $21.00 per hour. If 2,200 units used 22,900 hours at an hourly rate of $21.63 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.
Answer #1

Labor rate variance = (SR-AR) *AH

= (21 - 21.63)*22900

= 14,427 Unfavorable

Labor time variance = (SH-AH) * SR

= (2200*10 - 22900)*21

= 18,900 Unfavorable

Labor cost variance = 14,427U + 18900U

= 33,327 Unfavorable